TOPIC- SYSTEM OF EQUATION
how do you solve these two equations x+y=300 and 0.1x+3y=78 by using system of equations?

Answer by mananth(16946)

x+y=300.................(1)
0.1x+3y=78
multiply by 10
x+30y=780................(2)
1 x + 1 y = 300 .............1
Total value
1.00 x + 30.00 y = 780.00 .............2
Eliminate y
multiply (1)by -30.00
Multiply (2) by 1.00
-30.00 x -30.00 y = -9,000.00
1.00 x + 30.00 y = 780.00
Add the two equations
-29.00 x = -8,220.00
/ -29.00
x = 283.45
plug value of x in (1)
1.00 x + 1.00 y = 300.00
283.45 + y = 300.00
y = 300.00 -283
y = 16.55
y = 16.55